A long-distance runner runs the first kilometre of a race in 3 minutes 45 seconds but finds that his speed drops steadily such that each kilometre takes him 12 seconds more than the preceding one. Find his time (in minutes and seconds) for the 11th kilometre.


1
Expert's answer
2021-09-29T00:06:35-0400

If A1 = 225 seconds, and d = 12, then using the formula for arithmetic progression:

"A11 = A1 + (n - 1)\u00d7d = 225 + (11 - 1)\u00d712 = 345"

seconds or 5 minutes 45 seconds.